### v1

@@ Appendix A Extended Data
32
VinDr-CXR [49] Atelectasis, Lung tumor, Lung cavity, Tuberculosis, Pulmonary fibrosis, Clavicle fracture, Lung cyst, Pneumonia, Calcification, No finding, Rib fracture, Pleural thickening, Other disease, Mediastinal shift, Enlarged PA, Nodule/Mass, ILD, COPD, Pneumothorax, Consolidation, Infiltration, Pleural effusion, Other lesion, Cardiomegaly, Emphysema, Lung Opacity an image of {} 2
@@
32
WCE [78] normal, ulcerative colitis, polyps, esophagitis an endoscopic finding suggestive of {} 32
@@
32
Fundus JSIEC [81] Massive hard exudates, Vitreous particles, Possible glaucoma, Yellow-white spots-flecks, CRVO, BRVO, Large optic cup, Blur fundus without PDR, Bietti crystalline dystrophy, Rhegmatogenous RD, CSCR, Fibrosis, Congenital disc abnormality, Laser Spots, Vessel tortuosity, Maculopathy, RAO, Pathological myopia, DR2, Tessellated fundus, Retinitis pigmentosa, ERM, VKH disease, MH, Cotton-wool spots, Optic atrophy, Severe hypertensive retinopathy, DR3, DR1, Preretinal hemorrhage, Fundus neoplasm, Silicon oil in eye, Blur fundus with suspected PDR, Myelinated nerve fiber, Normal, Disc swelling and elevation, Peripheral retinal degeneration and break, Chorioretinal atrophy-coloboma, Dragged Disc {} presented in image 32
@@
8
VinDr-SpineXR [73] No finding, Osteophytes, Disc space narrowing, Surgical implant, Foraminal stenosis, Other lesions, Vertebral collapse, Spondylolysthesis a spine x-ray image of {} 2
NIH ChestX-ray14 [74] Atelectasis, Cardiomegaly, Effusion, Infiltration, Mass, Nodule, Pneumonia, Pneumothorax, Consolidation, Edema, Emphysema, Fibrosis, Pleural Thickening, Hernia, No Finding an image of {} 16
CheXpert [58] Enlarged Cardiomediastinum, Cardiomegaly, Lung Opacity, Lung Lesion, Edema, Consolidation, Pneumonia, Atelectasis, Pneumothorax, Pleural Effusion, Pleural Other, Fracture, Support Devices, No Finding an image of {} 4
ODIR [83] Normal, Diabetes, Glaucoma, Cataract, Age related Macular Degeneration, Hypertension, Pathological Myopia, Other diseases/abnormalities {} presented in retinography 32
BUID [84] normal, malignant, benign a close-up ultrasound scan showing a {} in the breast 32
@@
16
OCTMNIST [86] choroidal neovascularization, diabetic macular edema, drusen, normal an image of {} 4
Breast Tumor MRI [87] glioma, meningioma, pituitary, no tumor a breast mri image of {} 4
Retinal OCT [91] Age-related macular degeneration, Choroidal neovascularization, Central serous retinopathy, Diabetic macular edema, Macular hole, Drusen, Diabetic retinopathy, Normal a retinal oct image of {} 4
Brain Tumor CT [65] Healthy, Tumor an image of {} 4
COVIDxCT [92] Normal, Pneumonia, COVID-19 an image of {} 4
Brain Tumor MRI [65] Healthy, Tumor an image of {} 4
DDSM [66] negative, positive the mammography image of {} 4
Breast Cancer [69] Normal, Tumor the mammography image of {} 4
Table 3: AUC scores for classification results across different modalities in the zero-shot setting. Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
@@
Table 4: AUC scores for classification results across different modalities in the linear probes setting with 1% training data. Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
@@
Average 82.00 (81.01,82.84) 86.04 (85.52,86.55) 95.64 (95.16,96.01) 95.51 (95.06,95.96) 97.12 (96.78,97.42)
Table 5: AUC scores for classification results across different modalities in the linear probes setting with 10% training data. Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
@@
Average 91.76 (90.94,92.55) 94.98 (94.60,95.29) 97.70 (97.22,98.09) 96.99 (96.67,97.27) 98.40 (98.22,98.57)
Table 6: AUC scores for classification results across different modalities in the linear probes setting with 100% training data. Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
@@
Average 95.46 (94.79,96.03) 97.79 (97.56,98.01) 98.22 (97.82,98.56) 97.64 (97.40,97.84) 98.86 (98.71,98.99)
Table 7: AUC scores for classification results across different modalities in the fully fine-tuning setting. Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
@@
Table 8: Performance of Recall@1,5,10 metrics of different models on the PMC-9K dataset. Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
@@
Table 9: Performance of Recall@1,50,200 metrics of different models on the QUILT-1M dataset. Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
@@
Table 10: Results of models on the medical visual question answering task. Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
@@
Table 11: Results of models on the medical report generation task. Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
@@
BLEU-1 49.20 46.06 20.63 45.91 49.74
BLEU-2 32.32 29.51 9.90 29.10 32.57
BLEU-3 23.26 21.07 5.54 20.58 23.60
BLEU-4 17.37 15.74 3.19 15.44 17.82
ROUGE-L 37.64 37.40 16.98 35.95 39.06
METEOR 20.46 21.27 12.69 20.04 21.16
CIDEr 46.58 44.88 0.62 47.02 42.56
Micro Precision 54.70 55.71 19.52 51.36 51.84
Micro Recall 50.48 54.53 26.40 39.88 57.03
Micro F1 52.51 55.11 22.44 44.90 54.31
Table 12: Results of the models on various datasets for Whole-Slide image tasks. We use C-index for survival prediction tasks, for other tasks, we use AUC scores. Bold indicates the best result and underline indicates the second best. Mean±std is presented.
Dataset CLIP BiomedCLIP PMC-CLIP SigLIP-400M PLIP PathGen-CLIP ConceptCLIP
@@
BRACS-3 80.51±3.51 90.0±2.35 87.34±2.89 87.84±2.52 89.57±2.48 90.2±2.49 91.65±2.28
BRACS-7 72.68±3.27 81.43±2.68 77.29±3.07 80.85±2.78 82.57±2.54 84.97±2.11 85.04±2.18
BRCA 89.94±2.18 91.03±2.48 87.44±3.04 91.16±2.41 86.37±3.49 93.72±1.93 92.98±2.14
NSCLC 90.84±1.69 92.89±1.43 93.21±1.42 94.05±1.27 93.47±1.47 95.5±1.12 95.01±1.15
Camelyon 86.69±3.25 89.86±2.75 93.98±1.86 90.15±2.81 92.35±2.46 94.69±1.83 95.17±1.76
Molecular Subtyping
BRCA 70.23±2.93 72.15±2.96 71.75±3.09 70.92±2.64 73.93±3.05 74.04±2.81 74.36±3.08
@@
BRCA 59.16±5.72 62.5±5.76 57.22±5.78 62.49±5.15 59.45±5.75 66.35±4.41 64.59±5.24
LUAD 63.48±5.04 62.74±4.5 58.12±4.76 62.78±4.82 60.27±4.79 59.99±4.81 59.75±5.28
LUSC 58.4±4.25 55.98±4.63 55.5±4.27 56.86±4.82 54.83±4.43 56.88±4.43 59.66±4.77
Table 13: Performance of medical vision-language models on zero-shot medical concept annotation tasks (AUC %). “w/o Local Info.” denotes that the local information of image patches is not used in zero-shot concept annnotation. Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
Model Derm7pt (Dermoscopy) SkinCon (Dermoscopy) WBCAtt (Hematology) BrEaST (Ultrasound) LUNA16 (CT) Average
@@
Table 15: Ablation study on the effectiveness of the PC-Align loss and the local information (“Local Info.”). Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
@@
Table 16: Medical image analysis dataset splits and evaluation metrics.
Task Dataset Split Metrics
@@
Molecular Subtyping BRCA [94] $716$ / $102$ / $307$ AUC

### v2

@@ Appendix A Extended Data
VinDr-CXR [44] Atelectasis, Lung Tumor, Lung Cavity, Tuberculosis, Pulmonary Fibrosis, Clavicle Fracture, Lung Cyst, Pneumonia, Calcification, No Finding, Rib Fracture, Pleural Thickening, Other Disease, Mediastinal Shift, Enlarged PA, Nodule/Mass, ILD, COPD, Pneumothorax, Consolidation, Infiltration, Pleural Effusion, Other Lesion, Cardiomegaly, Emphysema, Lung Opacity an image of {}
@@
WCE [71] Normal, Ulcerative Colitis, Polyps, Esophagitis an endoscopic finding suggestive of {}
@@
Fundus JSIEC [74] Massive Hard Exudates, Vitreous Particles, Possible Glaucoma, Yellow-White Spots-Flecks, CRVO, BRVO, Large Optic Cup, Blur Fundus without PDR, Bietti Crystalline Dystrophy, Rhegmatogenous RD, CSCR, Fibrosis, Congenital Disc Abnormality, Laser Spots, Vessel Tortuosity, Maculopathy, RAO, Pathological Myopia, DR2, Tessellated Fundus, Retinitis Pigmentosa, ERM, VKH Disease, MH, Cotton-Wool Spots, Optic Atrophy, Severe Hypertensive Retinopathy, DR3, DR1, Preretinal Hemorrhage, Fundus Neoplasm, Silicon Oil in Eye, Blur Fundus with Suspected PDR, Myelinated Nerve Fiber, Normal, Disc Swelling and Elevation, Peripheral Retinal Degeneration and Break, Chorioretinal Atrophy-Coloboma, Dragged Disc {} presented in image
@@
VinDr-SpineXR [66] No Finding, Osteophytes, Disc Space Narrowing, Surgical Implant, Foraminal Stenosis, Other Lesions, Vertebral Collapse, Spondylolisthesis a spine x-ray image of {}
NIH ChestX-ray14 [67] Atelectasis, Cardiomegaly, Effusion, Infiltration, Mass, Nodule, Pneumonia, Pneumothorax, Consolidation, Edema, Emphysema, Fibrosis, Pleural Thickening, Hernia, No Finding an image of {}
CheXpert [50] Enlarged Cardiomediastinum, Cardiomegaly, Lung Opacity, Lung Lesion, Edema, Consolidation, Pneumonia, Atelectasis, Pneumothorax, Pleural Effusion, Pleural Other, Fracture, Support Devices, No Finding an image of {}
ODIR [76] Normal, Diabetes, Glaucoma, Cataract, Age Related Macular Degeneration, Hypertension, Pathological Myopia, Other Diseases/Abnormalities {} presented in retinography
BUID [77] Normal, Malignant, Benign a close-up ultrasound scan showing a {} in the breast
@@
OCTMNIST [79] Choroidal Neovascularization, Diabetic Macular Edema, Drusen, Normal an image of {}
Breast Tumor MRI [80] Glioma, Meningioma, Pituitary, No Tumor a breast mri image of {}
Retinal OCT [84] Age-Related Macular Degeneration, Choroidal Neovascularization, Central Serous Retinopathy, Diabetic Macular Edema, Macular Hole, Drusen, Diabetic Retinopathy, Normal a retinal oct image of {}
Brain Tumor CT [57] Healthy, Tumor an image of {}
COVIDxCT [85] Normal, Pneumonia, COVID-19 an image of {}
Brain Tumor MRI [57] Healthy, Tumor an image of {}
DDSM [58] Negative, Positive the mammography image of {}
Breast Cancer [61] Normal, Tumor the mammography image of {}
HMC-QU [62] Apical 4 Chamber, Apical 2 Chamber an image of {}
Figure 5: Concepts from the MedConcept-23M dataset. Wordclouds of concepts in captions to qualitatively visualize the concepts in each medical image modality. Word size is proportional to its occurrence in MedConcept-23M. Common nouns and verbs are ignored.
Table 3: AUC scores (%) for classification results across different modalities in the zero-shot setting. Bold indicates the best results and underline indicates the second best.The 95% CI is included in parentheses.
@@
Table 4: AUC scores (%) for classification results across different modalities in the linear probing setting with 1% training data. Bold indicates the best results and underline indicates the second best. The 95% CI is included in parentheses.
@@
Average 82.00 (69.56, 94.44) 86.04 (75.14, 96.95) 95.64 (92.98, 98.30) 95.51 (93.46, 97.57) 97.12 (95.31, 98.94)
Table 5: AUC scores (%) for classification results across different modalities in the linear probing setting with 10% training data. Bold indicates the best results and underline indicates the second best. The 95% CI is included in parentheses.
@@
Average 91.76 (88.09, 95.44) 94.98 (91.40, 98.55) 97.70 (96.82, 98.58) 96.99 (95.30, 98.68) 98.40 (97.35, 99.46)
Table 6: AUC scores (%) for classification results across different modalities in the linear probing setting with 100% training data. Bold indicates the best results and underline indicates the second best. The 95% CI is included in parentheses.
@@
Average 95.46 (93.87, 97.05) 97.79 (96.46, 99.13) 98.22 (97.59, 98.85) 97.64 (96.14, 99.13) 98.86 (98.11, 99.61)
Table 7: AUC scores (%) for classification results across different modalities in the fully fine-tuning setting. Bold indicates the best results and underline indicates the second best. The 95% CI is included in parentheses.
@@
Table 8: Recall@1, 5, 10 (%) of different models for cross-modal retrieval on the PMC-9K dataset. Bold indicates the best results and underline indicates the second best. The 95% CI is included in parentheses.
@@
Table 9: Recall@1, 50, 200 (%) of different models for cross-modal retrieval on the QUILT-1M dataset. Bold indicates the best results and underline indicates the second best. The 95% CI is included in parentheses.
@@
Table 10: Results (%) on the medical visual question answering task. Bold indicates the best results and underline indicates the second best. The 95% CI is included in parentheses.
@@
Table 11: Results (%) on the medical report generation task. Bold indicates the best results and underline indicates the second best. The 95% CI is included in parentheses.
@@
BLEU-1 44.11 (43.48, 45.64) 43.59 (42.59, 44.61) 20.31 (19.59, 21.00) 41.97 (40.89, 43.04) 44.58 (43.48, 45.64)
BLEU-2 27.95 (27.31, 29.82) 26.79 (25.64, 28.01) 8.75 (8.22, 9.30) 25.43 (24.23, 26.59) 28.60 (27.31, 29.82)
BLEU-3 19.08 (18.13, 20.66) 17.93 (16.86, 19.06) 3.97 (3.57, 4.37) 16.24 (14.94, 17.64) 19.38 (18.13, 20.66)
BLEU-4 12.57 (12.07, 14.51) 12.11 (11.09, 13.19) 1.61 (1.36, 1.87) 10.72 (9.59, 11.95) 13.30 (12.07, 14.51)
ROUGE-L 37.65 (37.39, 39.68) 37.37 (36.36, 38.45) 16.98 (16.39, 17.54) 35.97 (34.89, 37.06) 38.52 (37.39, 39.68)
METEOR 21.46 (20.81, 22.13) 22.06 (21.85, 23.31) 12.89 (12.42, 13.33) 21.14 (20.38, 21.93) 22.60 (21.85, 23.31)
Micro Precision 44.32 (40.17, 48.31) 47.74 (40.93, 48.81) 22.31 (18.94, 25.63) 35.01 (31.02, 38.81) 44.84 (40.93, 48.81)
Micro Recall 44.33 (40.17, 48.31) 47.24 (40.65, 48.36) 22.48 (19.24, 25.93) 34.75 (30.93, 38.39) 44.49 (40.65, 48.36)
Micro F1 44.26 (40.28, 48.36) 47.35 (40.68, 48.73) 22.29 (18.95, 25.62) 34.82 (31.02, 38.59) 44.52 (40.68, 48.73)
Table 12: Results (%) on various datasets for pathology whole-slide image tasks. We use the C-index for survival prediction tasks. For other tasks, we use AUC scores. Bold indicates the best results and underline indicates the second best. The 95% CI is included in parentheses.
Dataset CLIP BiomedCLIP PMC-CLIP SigLIP-400M ConceptCLIP
@@
BRACS-3 80.51 (73.63, 87.39) 90.00 (85.39, 94.61) 87.34 (81.68, 92.99) 87.84 (82.90, 92.78) 91.65 (87.18, 96.12)
BRACS-7 72.68 (66.27, 79.09) 81.43 (76.18, 86.68) 77.29 (70.27, 84.31) 80.85 (75.40, 86.30) 85.04 (80.77, 89.31)
NSCLC 90.84 (87.53, 94.15) 92.89 (89.09, 96.69) 93.21 (90.43, 96.00) 94.05 (91.56, 96.54) 95.01 (92.76, 97.26)
Camelyon 86.69 (80.32, 93.06) 89.86 (83.47, 96.25) 93.98 (90.33, 97.63) 90.15 (84.63, 95.67) 95.17 (91.72, 98.62)
Mutation Prediction
BRCA-TP53 60.83 (50.52, 71.14) 76.23 (67.21, 85.25) 66.89 (56.89, 76.89) 77.95 (69.19, 86.71) 80.15 (72.05, 88.25)
BRCA-TTN 42.74 (29.76, 55.72) 39.75 (23.26, 56.24) 40.03 (24.98, 55.08) 42.35 (26.77, 57.93) 52.14 (35.18, 69.10)
LUAD-TP53 45.61 (33.46, 57.76) 66.42 (55.35, 77.49) 60.96 (48.67, 73.25) 58.55 (46.69, 70.41) 68.02 (56.85, 79.19)
@@
BRCA 59.16 (47.95, 70.37) 62.50 (51.21, 73.79) 57.22 (46.00, 68.44) 62.49 (52.40, 72.58) 64.59 (54.32, 74.86)
LUSC 58.40 (50.07, 66.73) 55.98 (46.90, 65.06) 55.50 (47.13, 63.87) 56.86 (47.40, 66.32) 59.66 (50.31, 69.01)
Table 13: Performance of medical vision-language models on zero-shot medical concept annotation tasks (AUC %). “w/o Local Info.” denotes that the local information of image regions is not used in zero-shot concept annnotation. Bold indicates the best result and underline indicates the second best. 95% CI is included in parentheses.
Model Derm7pt (Dermatology) SkinCon (Dermatology) WBCAtt (Hematology) BrEaST (Ultrasound) LUNA16 (CT) Average
@@
Table 15: Ablation study on the effectiveness of the RC-Align loss and the local information (“Local Info.”) using the AUC (%) score. Bold indicates the best results and underline indicates the second best. The 95% CI is included in parentheses.
@@
Table 16: Medical image analysis dataset splits (training/validation/testing) and evaluation metrics.
Task Dataset Train/Valid/Test Split Metrics
@@
Mutation Prediction BRCA-TP53 [87] $438$ / $68$ / $128$ AUC
BRCA-TTN [87] $438$ / $68$ / $128$
LUAD-TP53 [87] $313$ / $44$ / $91$
@@
Table 17: Medical concept annotation datasets.
Dataset Description and setting Link
Derm7pt [12] A dermoscopic image dataset containing 1,011 images with clinical concepts for melanoma skin lesions in dermatology. The test set is adopted for evaluation. Only the dermoscopic images are considered in this paper for concept annotations. The seven clinical concepts include “Pigment network”, “Dots and globules”, “Pigmentation”, “Streaks”, “Regression structures”, “Blue-whitish veil”, and “Vascular structures”.
SkinCon [10] A skin disease dataset densely annotated by experts for fine-grained model debugging and analysis. We choose 22 concepts that have at least 50 images representing the concept from the Fitzpatrick 17k (F17k) [89] subset. 350 images are sampled for evaluation. The clinical concepts include “Papule”, “Plaque”, “Pustule”, “Bulla”, “Patch”, “Nodule”, “Ulcer”, “Crust”, “Erosion”, “Atrophy”, “Exudate”, “Telangiectasia”, “Scale”, “Scar”, “Friable”, “Dome-shaped”, “Brown (Hyperpigmentation)”, “White (Hypopigmentation)”, “Purple”, “Yellow”, “Black”, and “Erythema”.
WBCAtt [11] This dataset is a densely annotated dataset for white blood cell (WBC) recognition, containing 11 morphological attributes for 10,298 cell images. The test set which contains 3,099 images is used for evaluation. There are 11 concepts, including “Cell size”, “Cell shape”, “Nucleus shape”, “Nuclear cytoplasmic ratio”,“Chromatin density”, “Cytoplasm vacuole”, “Cytoplasm texture”, “Cytoplasm colour”, “Granule type”, “Granule colour”, and “Granularity”.
BrEaST [90] A breast ultrasound dataset containing 256 images with fine-grained concepts for breast lesion analysis. Seven BI-RADS descriptors are adopted as the evaluated concepts, including “Shape”, “Magrin”, “Echogenicity”, “Posterior features”, “Halo”, “Calcifications”, and “Skin thickening”.
LUNA16 [13] A curated version of LIDC-IDRI [91] dataset, which contains lung images with clinical concepts. 1,185 slices of lesions from the chest CTs with six clinical concepts are used for evaluation. The concepts include “Calcification”, “Lobulation”, “Margin”, “Sphericity”, “Spiculation”, and “Texture”.
Table 18: State-of-the-art vision-language pre-training models.
Model Name Description Link
Models in the General Domain
CLIP [34] CLIP [34] is a vision-language model pre-trained on large-scale image-text pairs via contrastive learning in the general domain, which showcases strong ability in zero-shot classification. We take the version of CLIP-ViT-B-16 pre-trained on LAION-2B [92] under the framework of OpenCLIP [93].
SigLIP-400M [28] SigLIP [28] is a variant of the CLIP model, which is pre-trained via contrastive learning using the sigmoid loss. In our experiments, we use the version of ViT-SO400M-14-SigLIP (SigLIP-400M).
Models in the Medical Domain
PMC-CLIP [3] PMC-CLIP [3] is a vision-language model tailored for medicine. The authors collected PubMed open access articles and conducted a series of data processing operations, producing a dataset with 1.6 million medical image-text pairs. The dataset is used to pre-train a CLIP model, thus resulting in the PMC-CLIP model. We use the origin checkpoint that the authors provided.
BiomedCLIP [2] BiomedCLIP [2] is a medical vision-language model pre-trained on PMC-15M, a dataset with 15 million medical image-text pairs sourced from PMC-OA. We take the released checkpoint for evaluation.